OpenStudy (anonymous):

Write an equation of the line containing the given point and parallel to the given line. (9,-7); 7x-4y=5 the equation of the line is y=?

OpenStudy (acxbox22):

first find slope of the given equation

OpenStudy (anonymous):

okay so how would i find the slope

OpenStudy (acxbox22):

7x-4y=5 -4y=-7x+5 divide by -4 y=7/4x-5/4 slope is 7/4

OpenStudy (anonymous):

okay

OpenStudy (acxbox22):

a parallel line is any line with the same slope to contain the given point you have to use point slope form

OpenStudy (acxbox22):

y+7=7/4(x-9) 7/4 is the slope so we keep that and we include point (9,-7) now you just have to change that into slope intercept form

OpenStudy (acxbox22):

y+7=7/4(x-9) y+7=7/4x-63/4 y=7/4x-91/4

OpenStudy (anonymous):

so the last one is the answer to the equation then ? you explain more well than most most people dont explain it good

OpenStudy (acxbox22):

y=7/4x-91/4

OpenStudy (anonymous):

thanks

OpenStudy (acxbox22):

sometimes i do these pretty fast so i am sorry if i made a mistake

OpenStudy (anonymous):

you didnt thanks though your a big help

OpenStudy (acxbox22):

np